OpenStudy (anonymous):

[6.02] Solve the following system. y = x + 3 2x + y = 9 (2, 5) (5, 2) (−2, 5) (2, −5)

OpenStudy (anonymous):

Do you know how to solve simultaneously?

OpenStudy (anonymous):

i'm just very confused with this whole lesson

OpenStudy (anonymous):

Okay - ill go through it with you So, you have ben given two equations.

OpenStudy (anonymous):

We want to know what x is so that it satisfies both equations.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

To do this, there is a method of subtitution.

OpenStudy (anonymous):

You have been given the equation (y= x+3) You have two unknows - x and y

OpenStudy (anonymous):

The first step is to substitute (y= x+3) into the the equation 2x+y=9 So you get: 2x+y =9 2x+(x+3)=9 Are you able to solve this?

OpenStudy (anonymous):

Do I solve the second one only?

OpenStudy (anonymous):

Yes, solve the second one

OpenStudy (anonymous):

HINT: 2x + (x+3) = 9 3x +3 = 9

OpenStudy (anonymous):